# Adjusting Model Size in Continual Gaussian Processes: How Big is Big Enough?

### v4

@@ 4 Automatically Adapting Approximation Capacity
We propose a method for adjusting the capacity of the approximation $M_{\bf b}$ to maintain accuracy. We propose to keep inducing points from old batches fixed, and select new inducing points from each incoming batch, with their locations set using the “greedy variance” criterion (Fine & Scheinberg, 2001; Foster et al., 2009; Burt et al., 2020). While optimising all inducing points does lead to a strictly better approximation, we avoid this for the sake of simplicity. The question remains: To achieve a level of accuracy, “how big is big enough?” To answer this, we consider the online ELBO as a function of the capacity ^$\widehat{\mathcal{L}}(M_{\bf b})$, and propose a threshold after which to stop adding new inducing variables.
@@ 4.1 Online Log Marginal Likelihood (LML) Upper Bound
The problem of selecting a sufficient number of inducing variables is also still open in the batch setting. One possible strategy is to use an upper bound to the marginal likelihood (Titsias, 2014) to bound $\mathrm{KL}[q(f)||p(f|\mathbf{y})]\leq\mathcal{U}-\mathcal{L}$, and stop adding inducing variables once this is below a tolerance $\alpha$. To extend this strategy to online learning, we begin by deriving an online upper bound, as a counterpart to the online ELBO from eq. (8). We follow the same strategy as Titsias (2014), by considering the highest possible value that our lower bound can attain. While in full-batch inference this is equal to the true LML, in our case this is obtained by keeping the inducing inputs from the previous iteration, and adding each new datapoint to the inducing set:
[…]
@@
$\leq-\frac{(N+M_{\mathbf{a}})}{2}\log(2\pi)-\frac{1}{2}\log|\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+\Sigma_{\mathbf{\hat{y}}}|$
@@
where $t=\operatorname{tr}(\mathbf{K}_{\mathbf{\hat{f}\hat{f}}}-\mathbf{Q}_{\mathbf{\hat{f}\hat{f}}})$ and $\mathbf{Q}_{\mathbf{\hat{f}\hat{f}}}=\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}$ and $M$ is the number of inducing points used to calculate the bound (which can be unequal to $M_{\mathbf{b}}$).
@@
  while $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\delta|\widehat{\mathcal{U}}(M)-\mathcal{L}_{noise}(\hat{\mu},\hat{\sigma})|$ do
@@ 4.2 Approximation Quality Guarantees
Adding inducing points will eventually increase $\widehat{\mathcal{L}}$ until it reaches $\mathcal{L}^{*}$ (Bauer et al., 2016; Matthews, 2017; Burt et al., 2020). If we add inducing points until $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\alpha$ we can guarantee the following:
@@
where $\Psi=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})}\mathrm{d}f$ and $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}\mid f)$ represents the variational distribution associated with the optimal lower bound =^$\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\mathbf{a}})$, with $\mathcal{Z}$ denoting the marginal likelihood that normalises $q_{n}^{*}(f)$.
@@
We cease the addition of points when ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})>\widehat{\mathcal{U}}(M)-\alpha$. Given that $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, and assuming $\theta_{n}=\theta_{o}$, the rest follows from algebraic manipulation of eq. (4). See App. C for the complete proof. ∎
@@ 4.3 Selecting a Threshold
In this final step of our online learning method, we must specify a heuristic for selecting $\alpha$ that does not require knowing any data in advance, while also working in a uniform way across datasets with different properties. A constant value for $\alpha$ does not work well, since the scale of the LML depends strongly on properties such as dataset size, and observation noise. This means that a tolerance of 1 nat (Cover, 1999) may be appropriate for a small dataset, but not for a large one.
As a principle for selecting the threshold, we take loose inspiration from compression and MDL (Grünwald & Roos, 2019), which takes the view of the ELBO being proportional to negative the code length that the model requires to encode the dataset. Intuitively, our desire to select an $\alpha$ such that our method captures a high proportion (e.g. 95%) of all the information in each batch, so that we can compress to within a small fraction of the optimal variational code. By itself, this is not well-defined since the code length depends on the quantisation tolerance for our continuous variables. To avoid this, we take an independent random noise code as our baseline, and select $\alpha$ to be within some small fraction of the optimal variational code, relative to the random noise code. We want to capture a high proportion of all the additional information that our model provides, relative to the noise model, i.e. we want our threshold to be:
@@
The procedure for our inducing point selection method is detailed in Algorithm 1 and we give further details in App. B.1. We name our approach Vegas Inducing Point Selection (VIPS), drawing an analogy to Las Vegas Algorithms. These methods guarantee the accuracy of the output, however, their computational time fluctuates for every run (Motwani & Raghavan, 1995).
@@ 5.2 The impact of model capacity on accuracy and training time
Here, we compare a fixed to a dynamic model size, with inducing points selected according to the greedy variance criterion (Burt et al., 2020), to demonstrate the challenges of a fixed model size. Since the total dataset size and data characteristics are not known at the start of training, selecting a fixed size will either lead to a model that is too small and underperforms, or a model that is too large and wastes computational resources. Adaptive size solves these issues.
[…]
@@ 5.3 Continual learning of UCI datasets
We compare our method, VIPS, to two other inducing point selection approaches: Conditional Variance (CV) and OIPS (Galy-Fajou & Opper, 2021) (see details in App. B). We use six datasets from the UCI repository (Dua & Graff, 2017), simulating a continual learning scenario by sorting the data along the first dimension and dividing it into batches. When evaluating inducing point selection methods, we must consider the trade-off between accuracy and model size: while selecting more inducing points improves model accuracy, it also increases computational cost.
@@
Table 1 shows the number of inducing points for each method at the operating point that achieves the accuracy threshold within 10% RMSE of a full-batch GP. VIPS requires fewer inducing points in the majority of datasets. CV often results in larger models, reaching the 7000-point limit on Bike. OIPS performs inconsistently, matching VIPS on smaller datasets but oversizing on larger ones. For the Naval dataset, OIPS failed to meet the accuracy constraint across all tested hyperparameter values. While extending the hyperparameter range might have improved accuracy, it would have led to even larger models. Therefore, we selected an operating point that satisfied our conditions for five out of six datasets, treating Naval as an outlier. Table 2 shows the results when this exception is applied to all methods. We observe a similar trend, with VIPS achieving the smallest model sizes. Tables 3 and 4 report analogous results for the threshold within 10% NLPD of a full-batch GP. The same pattern holds. Figure 3 summarises these outcomes across outlier allowances, with VIPS more frequently achieving minimal model size. App. D.3 provides additional results under different thresholds.
@@ 5.4 Continual Learning of Magnetic Field
We apply the optimal hyperparameters found in the previous section, selected under the 10% RMSE threshold, to the experimental setup from Chang et al. (2023a) using real-world data from Solin et al. (2018), where a robot maps magnetic field anomalies in a 6m x 6m indoor space. This scenario simulates an expanding domain, where the model needs to continuously learn new areas without uncontrollably increasing its capacity. Our goal is to test whether the methods can form representations by progressively spreading inducing points and learning hyperparameters without prior tuning in two scenarios: 1) sequentially mapping multiple robot trajectories and 2) processing a single path in continuous batches. Figure 4 shows the final magnetic field estimates produced by VIPS in both experimental settings. Comparative results for CV and OIPS are provided in App. D.4. The results show that VIPS was able to grow its memory as the robot moved, allowing for accurate modelling of the path and increasing its capacity dynamically. In contrast, we observe that CV increases its capacity uncontrollably while OIPS provides a less accurate estimate of the magnetic field due to choosing fewer inducing points.
@@ Acknowledgements
GPB is supported by EPSRC through the Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1.
@@ Appendix A Code
The methods discussed in this work, along with the code to reproduce our results, are available online at https://github.com/guiomarpescador/vips. All experiments and figures were done in Python using libraries TensorFlow v.2.12 (Abadi et al., 2016) and GPflow v.2.9 (Matthews et al., 2017). Both packages are under Apache License 2.0. The implementation of the online lower bound (see eq. (8)) by Bui et al. (2017) is available on GitHub under Apache license 2.0. For the “greedy variance” selection criterion and algorithm, we use the implementation provided in Burt et al. (2020) available under Apache license 2.0.
@@ Appendix B Adaptive Inducing Points Selection Methods
In sections 5.3 and 5.4 we compare our method, VIPS, to two other adaptive approaches: Conditional Variance (CV) and OIPS (Galy-Fajou & Opper, 2021). The key distinction between these approaches lies in their selection strategy of inducing points. OIPS uses a unified selection process (Algorithm 2), where the choice of both location and number of inducing points is determined by a single metric. In particular, each point $\mathbf{x}_{i}$ in the training set $X_{n}$ is evaluated individually, and those that meet the threshold condition are added to the inducing set $\mathbf{Z}_{n}$. In contrast, both CV and VIPS use a decoupled selection process (Algorithm 3), which separates two decisions: where to place inducing points and how many to include. These methods iteratively select the best candidate for the next inducing point from a pool of potential locations, continuing until a stopping criterion is met, allowing for more control over the model size. VIPS and CV share the same greedy variance selection criterion but differ in their candidate pools and stopping criteria. Table 5 summarises the properties of these methods, including the fixed memory approach of Bui et al. (2017) for comparison. The next three subsections provide detailed descriptions of each method’s specific implementation.
@@
  while stopping_criterion($\mathbf{Z}_{n}$) $>$ threshold do
@@
Bui et al. (2017) Fixed $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ Gradient optimisation M constant
OIPS (Galy-Fajou & Opper, 2021) Adaptive $\{\mathbf{X}_{n}\}$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\rho$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\bar{\rho}$
@@ B.1 Vegas Inducing Point Selection (VIPS) Algorithm
To select the location of our new inducing points we use the location selection strategy “greedy variance” proposed in Burt et al. (2020). This strategy iteratively selects points from a set based on a preference criterion until a stopping condition is met. In particular, it chooses the location of the next inducing point to maximise the marginal variance in the conditional prior $p(f_{\neq\mathbf{u}}|\mathbf{u})$. This is equivalent to maximising $\operatorname{diag}[\mathbf{K_{ff}}-\mathbf{Q_{ff}}]$. In continual learning, Chang et al. (2023a) use the “greedy variance” criterion by defining $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ as the selection pool from which inducing point locations are selected and maintaining a fixed number of inducing points. Similarly, Maddox et al. (2021) extends the “greedy variance” criterion to heteroskedastic Gaussian likelihoods and also uses a fixed memory approach. In our case, we tested the location strategy with our stopping criterion using both $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ and $\{\mathbf{X}_{n}\}$ as candidates pool for the locations of the inducing points. We did not find a substantial difference between the methods and hence opted for the simpler version where we keep the old inducing point locations fixed and choose the new set of inducing points from among the locations in $\{\mathbf{X}_{n}\}$.
Algorithm 4 presents an inducing point selection method using our stopping criterion combined with the location selection strategy “greedy variance”. The method takes as input a value for the hyperparameter $\theta_{n}$. In practice, we will set $\theta_{n}=\theta_{o}$ to select the number of inducing points; the hyperparameter $\theta_{n}$ is inferred by optimising ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})$ once the inducing point locations $\mathbf{Z}_{n}$ has been chosen. In Algorithm 1, $\widehat{\mathcal{U}}(M)$ is used to calculate the stopping criterion. However, in practice, since for the continual learning settings we investigate $N_{n}$ is small enough, we will use $\widehat{\mathcal{U}}(M_{\mathbf{a}}+N_{n})=\mathcal{L}^{*}$. This value is calculated once at the beginning of the process. The algorithm’s complexity depends on the number of inducing points $M_{\mathbf{b}}$ used to compute the lower bound ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})$ at each iteration. The computational complexity for calculating $\widehat{\mathcal{L}}$ at each batch is $O(N_{n}M_{\mathbf{b}}^{2}+M_{\mathbf{b}}^{3})$, and the memory requirement is $O(M_{\mathbf{b}}^{2})$, where $M_{\mathbf{b}}$ represents the total number of inducing points in the current batch.
@@
  while $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\delta|\widehat{\mathcal{U}}(M)-\mathcal{L}_{noise}(\hat{\mu},\hat{\sigma})|$ do

### v5

@@ 4 Automatically Adapting Approximation Capacity
We propose a method for adjusting the capacity of the approximation $M_{\bf b}$ to maintain accuracy. We propose to keep inducing points from old batches fixed, and select new inducing points from each incoming batch, with their locations set using the “greedy variance” criterion (fine2001efficient; foster2009stable; Burt2020). While optimising all inducing points does lead to a strictly better approximation, we avoid this for the sake of simplicity. The question remains: To achieve a level of accuracy, “how big is big enough?” To answer this, we consider the online ELBO as a function of the capacity $\widehat{\mathcal{L}}(M_{\bf b})$, and propose a threshold after which to stop adding new inducing variables.
@@ 4.1 Online Log Marginal Likelihood (LML) Upper Bound
The problem of selecting a sufficient number of inducing variables is also still open in the batch setting. One possible strategy is to use an upper bound to the marginal likelihood (titsias2014) to bound $\mathrm{KL}[q(f)||p(f|\mathbf{y})]\leq\mathcal{U}$-$\mathcal{L}$, and stop adding inducing variables once this is below a tolerance $\alpha$. To extend this strategy to online learning, we begin by deriving an online upper bound, as a counterpart to the online ELBO from eq. (8). We follow the same strategy as titsias2014, by considering the highest possible value that our lower bound can attain. While in full-batch inference this is equal to the true LML, in our case this is obtained by keeping the inducing inputs from the previous iteration, and adding each new datapoint to the inducing set:
$\mathcal{L}^{*}:=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})=\log\mathcal{N}\left(\hat{\mathbf{y}};\,\mathbf{0}\,,\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}+\Sigma_{\hat{y}}\right)+\Delta_{\mathbf{a}}$ (9)
@@
[…]
@@
[…]
@@
[…]
@@ 4.2 Approximation Quality Guarantees
Adding inducing points will eventually increase $\widehat{\mathcal{L}}$ until it reaches $\mathcal{L}^{*}$ (bauer2016; matthews2017thesis; Burt2020). If we add inducing points until $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\alpha$ we can guarantee the following:
@@
where $\Psi=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})}\mathrm{d}f$ and $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}\mid f)$ represents the variational distribution associated with the optimal lower bound $\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\mathbf{a}})$, with $\mathcal{Z}$ denoting the marginal likelihood that normalises $q_{n}^{*}(f)$.
@@
We cease the addition of points when $\widehat{\mathcal{L}}(M_{\mathbf{b}})>\widehat{\mathcal{U}}(M)-\alpha$. Given that $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, and assuming $\theta_{n}=\theta_{o}$, the rest follows from algebraic manipulation of eq. (4). See App. C for the complete proof. ∎
@@ 4.3 Selecting a Threshold
In this final step of our online learning method, we must specify a heuristic for selecting $\alpha$ that does not require knowing any data in advance, while also working in a uniform way across datasets with different properties. A constant value for $\alpha$ does not work well, since the scale of the LML depends strongly on properties such as dataset size, and observation noise. This means that a tolerance of 1 nat (cover1999elementsnats) may be appropriate for a small dataset, but not for a large one.
As a principle for selecting the threshold, we take loose inspiration from compression and MDL (grunwald2019minimum), which takes the view of the ELBO being proportional to negative the code length that the model requires to encode the dataset. Intuitively, our desire to select an $\alpha$ such that our method captures a high proportion (e.g. 95%) of all the information in each batch, so that we can compress to within a small fraction of the optimal variational code. By itself, this is not well-defined since the code length depends on the quantisation tolerance for our continuous variables. To avoid this, we take an independent random noise code as our baseline, and select $\alpha$ to be within some small fraction of the optimal variational code, relative to the random noise code. We want to capture a high proportion of all the additional information that our model provides, relative to the noise model, i.e. we want our threshold to be:
@@
The procedure for our inducing point selection method is detailed in Algorithm˜1 and we give further details in App. B.1. We name our approach Vegas Inducing Point Selection (VIPS), drawing an analogy to Las Vegas Algorithms. These methods guarantee the accuracy of the output, however, their computational time fluctuates for every run (VegasBook).
@@ 5.2 The impact of model capacity on accuracy and training time
Here, we compare a fixed to dynamic model size, with inducing points selected according to the greedy variance criterion (Burt2020), to demonstrate the challenges of a fixed model size. Since the total dataset size and data characteristics are not known at the start of training, selecting a fixed size will either lead to a model that is too small and underperforms, or a model that is too large and wastes computational resources. Adaptive size solves these issues.
Performance comparison: We train on a synthetic dataset divided into four batches, using a fixed model size, and our adaptive stopping criterion, VIPS. We test three fixed sizes: M=10, 20M=10,\,20, and $30$ inducing points. We record the test root mean square error (RMSE) and compare to an exact GP model with access to all current training data as a benchmark. Figure 2(a) shows that fixed size models $M=10,20$ lose accuracy with more data whereas fixed model M=30M\!=\!30 and VIPS match exact GP performance. At this point, we can select the model with $M=30$ with the benefit of hindsight. However, at the start of training, we could not have known that this size would have been sufficient at the point of testing. VIPS not only automatically ends up with this optimal size, but also avoids computational waste by using fewer inducing points in earlier batches.
@@ 5.3 Continual learning of UCI datasets
We compare our method, VIPS, to two other inducing point selection approaches: Conditional Variance (CV) and OIPS (Galy-Fajou2021-tu) (see details in App. B). We use six datasets from the UCI repository (UCI), simulating a continual learning scenario by sorting the data along the first dimension and dividing it into batches. When evaluating inducing point selection methods, we must consider the trade-off between accuracy and model size: while selecting more inducing points improves model accuracy, it also increases computational cost.
@@
Table 3: Mean (std) of inducing points in final batch across training/test splits at operating point selected to achieve 10% NLPD threshold. Cross (✗) indicates failure to meet accuracy threshold and "Max." denotes reaching maximum capacity limit. Values for OIPS are in italics, as the operating point was selected based on only five datasets, making it less stringent than other methods.
@@
Table 1 shows the number of inducing points for each method at the operating point that achieves the accuracy threshold within 10% RMSE of a full-batch GP. VIPS requires fewer inducing points in the majority of datasets. CV often results in larger models, reaching the 7000-point limit on Bike. OIPS performs inconsistently, matching VIPS on smaller datasets but oversizing on larger ones. For the Naval dataset, OIPS failed to meet the accuracy constraint across all tested hyperparameter values. While extending the hyperparameter range might have improved accuracy, it would have led to even larger models. Therefore, we selected an operating point that satisfied our conditions for five out of six datasets, treating Naval as an outlier. Table 2 shows the results when this exception is applied to all methods. We observe a similar trend, with VIPS achieving the smallest model sizes. Tables 3 and 4 report analogous results for the threshold within 10% NLPD of a full-batch GP. The same pattern holds. Figure˜3 summarises these outcomes across outlier allowances, with VIPS more frequently achieving minimal model size. App. D.3 provides additional results under different thresholds.
@@ 5.4 Continual Learning of Magnetic Field
We apply the optimal hyperparameters found in the previous section, selected under the 10% RMSE threshold, to the experimental setup from Chang2023 using real-world data from Solin2018, where a robot maps magnetic field anomalies in a 6m x 6m indoor space. This scenario simulates an expanding domain, where the model needs to continuously learn new areas without uncontrollably increasing its capacity. Our goal is to test whether the methods can form representations by progressively spreading inducing points and learning hyperparameters without prior tuning in two scenarios: 1) sequentially mapping multiple robot trajectories and 2) processing a single path in continuous batches. Figure˜4 shows the final magnetic field estimates produced by VIPS in both experimental settings. Comparative results for CV and OIPS are provided in App. D.4. The results show that VIPS was able to grow its memory as the robot moved, allowing for accurate modelling of the path and increasing its capacity dynamically. In contrast, we observe that CV increases its capacity uncontrollably while OIPS provides a less accurate estimate of the magnetic field due to choosing fewer inducing points.
@@ Acknowledgements
GPB is supported by EPSRC through the Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1. GPB acknowledges the support of G-Research for her participation in the conference.
@@ Appendix A Code
The methods discussed in this work, along with the code to reproduce our results, are available online at https://github.com/guiomarpescador/vips. All experiments and figures were done in Python using libraries TensorFlow v.2.12 (tensorflow) and GPflow v.2.9 (GPflow2017). Both packages are under Apache License 2.0. The implementation of the online lower bound (see eq. (8)) by Bui2017 is available on GitHub under Apache license 2.0. For the “greedy variance” selection criterion and algorithm, we use the implementation provided in Burt2020 available under Apache license 2.0.
@@ Appendix B Adaptive Inducing Points Selection Methods
In sections 5.3 and 5.4 we compare our method, VIPS, to two other adaptive approaches: Conditional Variance (CV) and OIPS (Galy-Fajou2021-tu). The key distinction between these approaches lies in their selection strategy of inducing points. OIPS uses a unified selection process (Algorithm 2), where the choice of both location and number of inducing points is determined by a single metric. In particular, each point $\mathbf{x}_{i}$ in the training set $X_{n}$ is evaluated individually, and those that meet the threshold condition are added to the inducing set $\mathbf{Z}_{n}$. In contrast, both CV and VIPS use a decoupled selection process (Algorithm 3), which separates two decisions: where to place inducing points and how many to include. These methods iteratively select the best candidate for the next inducing point from a pool of potential locations, continuing until a stopping criterion is met, allowing for more control over the model size. VIPS and CV share the same greedy variance selection criterion but differ in their candidate pools and stopping criteria. Table 5 summarises the properties of these methods, including the fixed memory approach of Bui2017 for comparison. The next three subsections provide detailed descriptions of each method’s specific implementation.
@@
while stopping_criterion($\mathbf{Z}_{n}$) >> threshold do
@@
Bui2017 Fixed $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ Gradient optimisation M constant
OIPS (Galy-Fajou2021-tu) Adaptive $\{\mathbf{X}_{n}\}$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\rho$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\bar{\rho}$
@@ B.1 Vegas Inducing Point Selection (VIPS) Algorithm
To select the location of our new inducing points we use the location selection strategy “greedy variance” proposed in Burt2020. This strategy iteratively selects points from a set based on a preference criterion until a stopping condition is met. In particular, it chooses the location of the next inducing point to maximise the marginal variance in the conditional prior $p(f_{\neq\mathbf{u}}|\mathbf{u})$. This is equivalent to maximising $\operatorname{diag}[\mathbf{K_{ff}}-\mathbf{Q_{ff}}]$. In continual learning, Chang2023 use the “greedy variance” criterion by defining $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ as the selection pool from which inducing point locations are selected and maintaining a fixed number of inducing points. Similarly, Maddox extends the “greedy variance” criterion to heteroskedastic Gaussian likelihoods and also uses a fixed memory approach. In our case, we tested the location strategy with our stopping criterion using both $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ and $\{\mathbf{X}_{n}\}$ as candidates pool for the locations of the inducing points. We did not find a substantial difference between the methods and hence opted for the simpler version where we keep the old inducing point locations fixed and choose the new set of inducing points from among the locations in $\{\mathbf{X}_{n}\}$.
Algorithm 4 presents an inducing point selection method using our stopping criterion combined with the location selection strategy “greedy variance”. The method takes as input a value for the hyperparameter $\theta_{n}$. In practice, we will set $\theta_{n}=\theta_{o}$ to select the number of inducing points; the hyperparameter $\theta_{n}$ is inferred by optimising $\widehat{\mathcal{L}}(M_{\mathbf{b}})$ once the inducing point locations $\mathbf{Z}_{n}$ has been chosen. In Algorithm 1, $\widehat{\mathcal{U}}(M)$ is used to calculate the stopping criterion. However, in practice, since for the continual learning settings we investigate $N_{n}$ is small enough, we will use $\widehat{\mathcal{U}}(M_{\mathbf{a}}+N_{n})=\mathcal{L}^{*}$. This value is calculated once at the beginning of the process. The algorithm’s complexity depends on the number of inducing points $M_{\mathbf{b}}$ used to compute the lower bound $\widehat{\mathcal{L}}(M_{\mathbf{b}})$ at each iteration. The computational complexity for calculating $\widehat{\mathcal{L}}$ at each batch is $O(N_{n}M_{\mathbf{b}}^{2}+M_{\mathbf{b}}^{3})$, and the memory requirement is $O(M_{\mathbf{b}}^{2})$, where $M_{\mathbf{b}}$ represents the total number of inducing points in the current batch.
@@
[…]